Q1. Consider a steady, two-dimensional, incompressible flow field has the velocity potential a) b) c) 2 (x-7)(x+y) Determinecan you solve the last question (e)
Answer #1

e) A two dimensional, steady in compressible velocity flow es given by ū=ui + vj where u = ax + b V = - ay tcx The two dimensx component (Pe). an s [(ax+b) a + (- ay tez):o) : evo + 2 (0+0] -alx s [a²x + ab لا ax Pa - sha²x + ab] de Pa = -8 [a²x² + aa² : Py = -S Lary + bcy ] + f(a) + but pro,o=0 11 - loto] + f(x) + (2 f(x) + z = 0 b c I Py=-9 + cy 2 components of pressure